# Probing supermassive black hole scalarization with Pulsar Timing Arrays

### v1

@@ V Conclusions
For this reason, PTAs stand out as uniquely suited experiments to test this theory. PTAs monitor a network of millisecond pulsars over decade-long experiment. The recent detection of a temporally-correlated common stochastic noise, featuring a distinctive correlation among pulsars TOAs, has opened a window not only into astrophysical phenomena, but also potential signature of new physics (see e.g. Refs [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 27, 28, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95] and references therein). In this work, we assume that the recently observed common excess in pulsars’ TOAs arises from the incoherent superposition of GW emitted by circularly inspiralling SMBHBs. We discuss how the specific variant of EsGB gravity from Ref. [18] and described in Sec. II impacts the characteristic strain generated by the underlying astrophysical population. These effects are illustrated in Fig. 2 , which shows that this model leads to a flattening of the spectral index compared to the prediction from circular binaries in pure GR. Notably, this behavior aligns with current PTA observations, as shown in Fig. 3 for the EPTA DR2New dataset. Furthermore, as the EsGB theory of gravity predicts a regime dominated by dipolar emission for binaries involving at least one scalarized SMBH, we also observe a decrease in the overall amplitude of the characteristic strain with respect to GR, keeping the parameters in the merger density fixed. It may be interesting to confront this prediction with astrophysically-motivated priors on the merger density parameters, as done in Refs. [27, 96]. In particular, EsGB gravity may shift the amplitude toward the high end of the prior range, leading to a slight tension. We leave this aspect for future work. After analyzing the main features of EsGB gravity using mock data in Sec. IV.1, we consider real PTA observations in Sec. IV.2. In the main text, we report results for the EPTA DR2New dataset, where we find a marginal preference for values of $\lambda$ in the range $1.5\cdot 10^{7}~{}\text{km}\lesssim\lambda\lesssim 1.5\cdot 10^{9}~{}\text{km}$. However, it is important to emphasize that several astrophysical mechanisms, as high eccentricities or interaction with the environment [31, 32, 33, 34, 35, 36, 37], could also lead to a flattening of the spectral index, making them largely degenerate with the signatures expected from EsGB gravity. Future data releases, with longer observation times and enhanced precision, will definitely advance our comprehension of the astrophysical processes influencing the SGWB produced by SMBHBs. These advancements may help disentangle the different effects and ultimately lead to a final conclusion.
@@ VI Acknowledgements
We warmly thank Alberto Sesana for insightful discussions and important suggestions. Moreover, we wish to thank Enrico Barausse and Aaron Held for helpful comments and discussions on the manuscript and Nataliya Porayko for kindly providing a code to reproduce the free spectrum of timing residuals for the EPTA dataset. L.C. acknowledges support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673), the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” (Grant No. GWSky-101167314), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”, and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. A.K. acknowledges funding from the FCT project “Gravitational waves as a new probe of fundamental physics and astrophysics” grant agreement 2023.07357.CEECIND/CP2830/CT0003.
Appendix A Binary dynamics in the post-Newtonian approximation
In this section we want to provide a schematic picture of the derivation of the equations for the evolution of the orbital frequency that we employed in the main text. Here we summarize the main steps; for a more in-depth discussion, the reader is referred to Ref. [97, 98].
@@ Appendix A Binary dynamics in the post-Newtonian approximation
$\begin{split}G_{\mu\nu}=&2\partial_{\mu}\varphi\partial_{\nu}\varphi-g_{\mu\nu}\left((\partial\varphi)^{2}-\frac{\lambda^{2}}{2}F(\varphi)f(\mathcal{G})\right)\\ &+4\lambda^{2}\vphantom{}{}^{**}R_{\mu}{}^{\alpha\beta}{}_{\nu}\nabla_{\alpha}\nabla_{\beta}\left(F(\varphi)f^{\prime}(\mathcal{G})\right)+\\ &+8\pi GT_{\mu\nu}^{A}+(A\leftrightarrow B)\,,\\ \Box\varphi=&\frac{\lambda^{2}}{4}F^{\prime}(\varphi)f(\mathcal{G})\\ &+4\pi G\int\mathrm{d}s_{A}\frac{{\rm d}m_{A}}{{\rm d}\varphi}\frac{\delta^{(4)}(x-x_{A}(s_{A}))}{\sqrt{-g}}\\ &+(A\leftrightarrow B)\,,\end{split}$ (37)
@@
$\begin{split}&g_{00}=\exp{\left(-2U\right)}+\mathcal{O}(v^{6})\,,\\ &g_{i0}=-4g_{i}+\mathcal{O}(v^{5})\,,\\ &g_{ij}=\delta_{ij}\exp{\left(2U\right)}+\mathcal{O}(v^{6})\,,\end{split}$ (40)
@@
$\begin{split}&\varphi=\varphi_{0}+\delta\varphi+\mathcal{O}(v^{6})\,,\\ &m_{A}(\varphi)=M_{A}\left(1-\alpha_{A}\delta\varphi+\left((\alpha_{A})^{2}-\alpha_{A}^{\prime}\right)\delta\varphi^{2}\right)\\ &\qquad\qquad+\mathcal{O}(v^{6})\,,\end{split}$ (41)
@@
$\begin{split}\frac{{\rm d}f_{s}}{{\rm d}t}\simeq&\frac{\,x^{9/2}}{\pi G_{\text{eff}}^{2}\mathcal{M}^{2}(1+\alpha_{A}\alpha_{B})}\left(\eta^{2/5}\Delta\alpha^{2}\right.\\ &\left.+\frac{4}{5}\left(24+(\alpha_{A}+\alpha_{B})^{2}\right)x\right)\,,\end{split}$ (51)
@@
Appendix B Supplemental plots
@@
In this section, we present some plots that integrate the discussion in the main text. In Fig. 7, we show the posteriors on $\log_{10}(A/\text{Mpc}^{-3})$ and $\log_{10}\lambda_{9}$ for the idealized analysis described in Sec. IV.1. The upper panel shows $\epsilon=1/50$, the lower panel shows $\epsilon=1/20$. As $\epsilon$ increase, the constraints become wider, signaling a precision loss on the recovery of the injected parameters (solid blue lines). However, the characteristic shape in the 2D corner plot remains, and can be considered as a distinctive signature of EsGB data. Fig. 8 presents the posterior distributions of the parameters describing the merger density model in Eq.(15), obtained from the EPTA DR2New dataset using the first four frequency bins. Finally, Fig. 9 shows the results for the NANOGrav 15yr dataset, using the first fourteen frequency bins.

### v2

@@ V Conclusions
For this reason, PTAs stand out as uniquely suited experiments to test this theory. PTAs monitor a network of millisecond pulsars over decade-long experiment. The recent detection of a temporally-correlated common stochastic noise, featuring a distinctive correlation among pulsars TOAs, has opened a window not only into astrophysical phenomena, but also potential signature of new physics (see e.g. Refs [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 27, 28, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99] and references therein).
In this work, we assume that the recently observed common excess in pulsars’ TOAs arises from the incoherent superposition of GW emitted by circularly inspiralling SMBHBs. We discuss how the specific variant of EsGB gravity from Ref. [18] and described in Sec. II impacts the characteristic strain generated by the underlying astrophysical population. These effects are illustrated in Fig. 2 , which shows that this model leads to a flattening of the spectral index compared to the prediction from circular binaries in pure GR. Notably, this behavior aligns with current PTA observations, as shown in Fig. 3 for the EPTA DR2New dataset. Furthermore, as the EsGB theory of gravity predicts a regime dominated by dipolar emission for binaries involving at least one scalarized SMBH, we also observe a decrease in the overall amplitude of the characteristic strain with respect to GR, keeping the parameters in the merger density fixed. It may be interesting to confront this prediction with astrophysically-motivated priors on the merger density parameters, as done in Refs. [27, 100]. In particular, EsGB gravity may shift the amplitude toward the high end of the prior range, leading to a slight tension. We leave this aspect for future work.
After analyzing the main features of EsGB gravity using mock data in Sec. IV.1, we consider real PTA observations in Sec. IV.2. In the main text, we report results for the EPTA DR2New dataset, where we find a marginal preference for values of $\lambda$ in the range $1.5\times 10^{7}~\text{km}\lesssim\lambda\lesssim 1.5\times 10^{9}~\text{km}$. However, it is important to emphasize that several astrophysical mechanisms, as high eccentricities or interaction with the environment [31, 32, 33, 34, 35, 36, 37], could also lead to a flattening of the spectral index, making them largely degenerate with the signatures expected from EsGB gravity. Future data releases, with longer observation times and enhanced precision, will definitely advance our comprehension of the astrophysical processes influencing the SGWB produced by SMBHBs. These advancements may help disentangle the different effects and ultimately lead to a final conclusion.
@@ VI Acknowledgements
We warmly thank Alberto Sesana for insightful discussions and important suggestions. Moreover, we wish to thank Enrico Barausse and Aaron Held for helpful comments and discussions on the manuscript, Nataliya Porayko for kindly providing a code to reproduce the free spectrum of timing residuals for the EPTA dataset and Rahul Srinivasan for useful discussions about Bayesian inference. L.C. acknowledges support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673), the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” (Grant No. GWSky-101167314), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”, and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. A.K. acknowledges funding from the FCT project “Gravitational waves as a new probe of fundamental physics and astrophysics” Grant Agreement No. 2023.07357.CEECIND/CP2830/CT0003.
Appendix A BINARY DINAMICS IN THE POST-NEWTONIAN APPROXIMATION
In this section, we want to provide a schematic picture of the derivation of the equations for the evolution of the orbital frequency that we employed in the main text. Here we summarize the main steps; for a more in-depth discussion, the reader is referred to Refs. [101, 102].
@@ Appendix A BINARY DINAMICS IN THE POST-NEWTONIAN APPROXIMATION
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix B BAYESIAN MODEL COMPARISON
In the main text, we reported a mild excess in the posterior of $\log_{10}\lambda_{9}$, suggesting a slight preference for the EsGB theory over GR in explaining PTA data. However, the definitive statistical preference for a model over another is quantified by the Bayes factor, defined as the ratio of the evidences $Z$ of the two models. In formulae:
$B_{\text{10}}=\frac{p(\log_{10}\delta t|\mathcal{M}_{1})}{p(\log_{10}\delta t|\mathcal{M}_{0})}\equiv\frac{Z_{1}}{Z_{0}},$ (54)
where $\mathcal{M}_{1}$ and $\mathcal{M}_{0}$ denote the EsGB and GR theories, respectively. As expressed in Eq. (54), the Bayes factor $B_{10}$ quantifies whether it is more probable that the PTA data, collected in $\log_{10}\delta t$, are generated from EsGB ($\mathcal{M}_{1}$) or from GR ($\mathcal{M}_{0}$). A value $B_{10}>1$ favors EsGB, $B_{10}<1$ favors GR, and $B_{10}=1$ indicates that the two models are equally likely.
The Bayes factor in Eq. (54) can be rewritten as
$B_{10}=\frac{\int\mathcal{L}(\log_{10}\delta t|\theta_{1},\mathcal{M}_{1})~\pi(\theta_{1}|\mathcal{M}_{1})~d\theta_{1}}{\int\mathcal{L}(\log_{10}\delta t|\theta_{0},\mathcal{M}_{0})~\pi(\theta_{0}|\mathcal{M}_{0})~d\theta_{0}},$ (55)
where the evidences are expressed as the product of the likelihoods and the priors, integrated over the model parameters $\theta_{1}$ and $\theta_{0}$. Incidentally, this also highlights why the evidence is also referred to as marginalized likelihood. The likelihood for EsGB is given in Eq. (36), as a function of the parameters $\theta_{1}=\{\log_{10}\frac{\dot{n}_{0}}{\text{Mpc}^{-3}\text{Gyr}^{-1}},\log_{10}\lambda_{9},\beta,z_{0},\gamma,\log_{10}\mathcal{M}_{\star}/\text{M}_{\odot}\}$. For GR, the likelihood has the same functional form, but with $\lambda_{9}=0$, resulting in the set of parameters $\theta_{0}=\{\log_{10}\frac{\dot{n}_{0}}{\text{Mpc}^{-3}\text{Gyr}^{-1}},\beta,z_{0},\gamma,\log_{10}\mathcal{M}_{\star}/\text{M}_{\odot}\}$. For simplicity, we will consider only the case where the mass ratio $q$ follows a log-uniform distribution over $\log_{10}q\in[-1.5,0]$; using a uniform distribution yields similar results.
The integrals in Eq. (55) are evaluated via Monte-Carlo integration. In practice, the evidences are approximated by drawing $N$ random realizations of the collection of parameters $\theta_{k}$ for the model $k=0,1$ and computing
$Z_{k}\sim\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\log_{10}\delta t|\theta_{ki},\mathcal{M}_{k}),$ (56)
where $\theta_{ki}$ is the $i$-th random draw of the parameters for the model $\mathcal{M}_{k}$. This approximation assumes that the priors on $\theta_{k}$ for the model $\mathcal{M}_{k}$ are normalized and uniform within the integration interval, which is verified in our analysis.
Given the estimation of $Z_{k}$ through Eq. (56), we compute the associated uncertainties as
$\sigma_{k}^{2}=\frac{1}{N}\sum_{i=1}^{N}\bigg[\mathcal{L}(\log_{10}\delta t|\theta_{ki},\mathcal{M}_{k})-Z_{k}\bigg]^{2}$ (57)
$\sigma_{Z_{k}}^{2}=\frac{1}{N}\sigma^{2}_{k}.$ (58)
Figure 7 illustrates this procedure for the EPTADR2New dataset with $N_{\text{bins}}=9$. The evidences for EsGB and GR are shown in solid blue and dashed orange, respectively, and are computed using Eqs. (56) and (58) for an increasing number of random samples. Convergence is reached for $N\gtrsim 10^{4}$, and the error bars follow the scaling predicted by Eq.(58).
Figure 7 clearly shows that the EsGB theory explains the data slightly better than GR, as the Bayes factor in Eq. (54) turns out to be larger than one. However, substituting the converged values in Eq. (54), we find $B_{10}\sim 1.35$ for a large number of samples, corresponding to a “barely worth mentioning” preference, according to Jeffreys’ [106] and Kass and Raftery’s [107] scales. In other words, even though the Bayes factor mildly favors EsGB over GR, the preference is largely inconclusive. Future data will be essential to support a more definitive conclusion. We performed the same analysis for the NANOGrav 15-year dataset and for the EPTADR2New with $N_{\text{bins}}=4$, obtaining similar results. The corresponding plots are omitted for brevity.
Figure 7: Monte-Carlo estimate of the evidence $Z$ with associated uncertainty $\sigma_{Z}$, as computed by Eqs. (56) and (58), for an increasing number of random samples. The solid blue line is computed assuming the EsGB theory of gravity, while the dashed orange line shows the evidence for GR. The plot is produced for the EPTADR2New dataset, using $N_{\text{bins}}=9$.
To further confirm our findings, we also compute the Bayes factor using the Savage-Dickey density ratio [108, 109, 110]. This method is applicable when one model is nested within the other - that is, when the former can be derived by setting certain parameters of the latter to fixed values. This is exactly our case: indeed, GR is obtained by setting $\lambda_{9}=0$ in EsGB theory or, equivalently, $\log_{10}\lambda_{9}=$-$\infty$. Formally, we can write $\theta_{0}=\theta_{1}|_{\log_{10}\lambda_{9}=-\infty}$.
The further requirement $\pi(\theta_{0}|\mathcal{M}_{0})=\pi(\theta_{0}|\log_{10}\lambda_{9}=-\infty,\mathcal{M}_{1})$ is trivially satisfied in our case, as the priors are fully separable. In this case, the Bayes factor reads:
$B_{10}=\frac{\pi(\log_{10}\lambda_{9}=-\infty|\mathcal{M}_{1})}{\mathcal{P}(\log_{10}\lambda_{9}=-\infty|\log_{10}\delta t,\mathcal{M}_{1})},$ (59)
and is given by the ratio of the prior and posterior probability densities at $\log_{10}\lambda_{9}=$-$\infty$, both evaluated in the EsGB model.
The advantage of this method lies in its efficiency: it avoids computing $Z_{0}$ and $Z_{1}$ separately, which can be computationally expensive, and requires only prior and posterior samples from the more general model $\mathcal{M}_{1}$.
The intuition behind this approach is that if the data favor EsGB over GR, the posterior density at $\log_{10}\lambda_{9}=$-$\infty$ becomes smaller than the prior density at that point, resulting in a Bayes factor greater than one.
In our analysis, the prior on $\log_{10}\lambda_{9}$ does not extend all the way to -$\infty$; however, values of $\log_{10}\lambda_{9}$ close to the lower end of the prior range are effectively indistinguishable from GR for our purposes.
Therefore, as a rough approximation, we evaluate the Savage-Dickey density ratio at $\log_{10}\lambda_{9}=-5$. For the EPTADR2New dataset with $N_{\text{bins}}=9$, this yields $B_{10}\sim 1.2$, which is broadly in agreement with the result obtained via Monte-Carlo direct integration. Accordingly, in the main text we only report Bayes factors obtained from direct evidence integration, noting that they are consistent with alternative estimation methods.
Appendix C SUPPLEMENTAL PLOTS
@@
In this section, we present some plots that integrate the discussion in the main text. In Fig. 8, we show the posteriors on $\log_{10}(A/\text{Mpc}^{-3})$ and $\log_{10}\lambda_{9}$ for the idealized analysis described in Sec. IV.1. The upper panel shows $\epsilon=1/50$, the lower panel shows $\epsilon=1/20$. As $\epsilon$ increase, the constraints become wider, signaling a precision loss on the recovery of the injected parameters (solid blue lines). However, the characteristic shape in the 2D corner plot remains, and can be considered as a distinctive signature of EsGB data. Figure 9 presents the posterior distributions of the parameters describing the merger density model in Eq.(15), obtained from the EPTA DR2New dataset using the first four frequency bins. Finally, Fig. 10 shows the results for the NANOGrav 15yr dataset, using the first fourteen frequency bins.